# MedChain: Bridging the Gap Between LLM Agents and Clinical Practice through Interactive Sequential Benchmarking

### v1

@@ 4.1.2 Decision Making
Single-agent GPT-4o-mini 2023 0.5449 0.2871 0.3399 0.5112 0.4406 0.3930 0.4195
Claude-3.5-sonnet 2023 0.5681 0.3050 0.3562 0.5018 0.4599 0.4053 0.4327
LLaVA 2024a 0.3240 0.0730 0.3182 0.5165 0.4024 0.1060 0.2901
Qwen2 2024a 0.4975 0.2215 0.4226 0.4829 0.4818 0.2193 0.3876
InternVL2 2024a 0.4811 0.1935 0.4645 0.4490 0.4738 0.2903 0.3920
Multi-agent MedAgent 2023 0.3830 0.2039 0.4454 0.4456 0.4266 0.3673 0.3786
MDAgent 2024a 0.2398 0.1343 0.4240 0.4983 0.4289 0.3620 0.3479
MedChain-Agent 0.5873 0.3505 0.5836 0.6566 0.5218 0.4613 0.5269
4.2 MedCase-RAG
To enhance the decision-making capabilities of our multi-agent framework, we introduce MedCase-RAG, a novel Retrieval-Augmented Generation technique tailored for medical contexts.
Unlike conventional medical RAG methods that rely on unstructured medical QA databases, MedCase-RAG employs a structured approach to data representation. We map each medical case into a 12-dimensional feature vector, encompassing crucial attributes such as Age, Sex, Patient Description, Symptom Description, Patient History, and Patient Image. This structured representation allows for more comprehensive and standardized case encoding, facilitating more accurate and nuanced retrieval. Within this feature set, we identify "Symptom Description" as the most representative feature of a patient’s condition. This feature undergoes quantification through a Text Embedding model, serving as the primary key for dense retrieval tasks. Moreover, MedCase-RAG incorporates a dynamic database expansion mechanism to continually enrich its knowledge base. As patients complete their medical processes and recover, their case information is reintroduced into the database as pseudo-data. This approach allows the system to learn from new cases and adapt to evolving medical knowledge and practices.
When a new patient case is presented, MedCaseRAG performs a similarity search within the relevant medical department, using cosine similarity to identify the top three most similar cases, offering a more comprehensive basis for decision-making, allowing agents to consider various aspects and potential outcomes.
@@ 5.1 Experimental Setup
We split the dataset into training, validation, and testing sets with a ratio of 7:1:2. For frameworks incorporating RAG techniques, both training and validation sets are used to construct the case retrieval database. Our study evaluates both single-agent and multi-agent systems. For single agents, we test two closed-source models (gpt-4o-mini OpenAI (2023), and claude-3.5-sonnet Team et al. (2023)) and three open-source models (InternVL2-8b Chen et al. (2023), llava-llama-3-8b-v1_1 Liu et al. (2024a), and Qwen2-7B-Instruct Wang et al. (2024a)), with model weights obtained from official Hugging Face repositories. In the multi-agent evaluation, we compare MedChain-Agent against MedAgent Tang et al. (2023) and MDAgent Kim et al. (2024a). All multi-agent framework are based on InternVL2-8b Chen et al. (2023). The deployment was conducted using the LMDeploy framework Contributors (2023). All tests executed on NVIDIA A100 GPUs featuring 80GB of memory. To enhance output stability and reliability across all experiments, we consistently set the temperature parameter to 0.
@@
0.5523 0.2228 0.3285 0.6369 0.4982 0.3915 0.4384
0.5739 0.2906 0.4222 0.6377 0.5086 0.4209 0.4757
0.5928 0.3353 0.5801 0.6488 0.5035 0.4568 0.5195
0.5873 0.3505 0.5836 0.6566 0.5218 0.4613 0.5269
5.2 Evaluation Metrics
To assess model performance, we employ diverse metrics tailored to each task. Accuracy is used for tasks with a single correct answer, such as Specialty Referral (Level 1). For tasks involving multiple key points, including Specialty Referral (Level 2), History-taking, and Treatment, we apply Intersection over Union (IoU) to capture the overlap between predicted and ground truth sets. Specifically for the Examination task, we utilize DocLens Xie et al. (2024) to evaluate the quality of image interpretation reports.
5.3 Experiment Result and Analysis
The results of our evaluation in the MedChain  are presented in Table 2. Our analysis yields three significant insights:
(1) Sequential decision-making tasks continue to pose significant challenges, even for advanced models. For instance, within the single-agent frameworks, GPT-4o-mini and InternVL2 achieve average scores of 0.4442 and 0.3920, respectively. These results indicate that despite their sophistication, these models struggle to maintain consistent performance across the sequential stages of clinical decision-making, highlighting the inherent difficulty of these tasks.
(2) Multi-agent frameworks based on InternVL2, such as MedAgent and MDAgent, exhibit inferior performance compared to their single-agent counterparts. This degradation suggests that traditional multi-agent approaches may exacerbate error propagation in sequential decision-making processes, leading to reduced overall performance. In contrast, our proposed MedChain-Agent significantly outperforms these multi-agent methods, achieving an average score of 0.5269. This improvement demonstrates that MedChain-Agent effectively mitigates error propagation, enhancing decision quality and reliability in clinical settings.
(3) The integration of the MedChain-Agent framework with open-source LLMs demonstrates significant superiority over proprietary models like GPT-4o-mini. The substantial performance gain observed with MedChain-Agent (average score of 0.5269) implies that our framework can leverage the strengths of open-source LLMs to achieve superior outcomes. This suggests that open-source models, when enhanced with our framework, are not only competitive but can also excel in handling intricate medical decision-making tasks.
5.4 Ablation Study
@@
Full MedAgent 0.4266 0.3673
MDAgent 0.4289 0.3620
gpt-4o-mini 0.4406 0.3930
InternVL2 0.4759 0.4472
w/o Person. $\downarrow$ MedAgent 0.3274 $\downarrow$ 0.3754 $\uparrow$
@@
[…]
w/o Inter. $\uparrow$ MedAgent 0.3315 $\downarrow$ 0.4109 $\uparrow$
MDAgent 0.4295 $\uparrow$ 0.3627 $\uparrow$
[…]
Three key characteristics for MedChain: To validate the effectiveness of personalization, interactivity, and sequentiality within our benchmark, we conduct ablation study as in Table 4. We systematically remove each characteristic and observe its impact on model performance across Diagnosis and Treatment tasks. 1) Removing patient-specific information (w/o Person.) consistently degrades performance across all models in diagnosis tasks (with drops ranging from 3.0% to 13.4%). This demonstrates that personalized information is crucial for accurate clinical decision-making. 2) Interestingly, when removing the sequential dependency between stages (w/o Seq.), most models show improved performance. This improvement actually validates the effectiveness of sequentiality in our benchmark, as it indicates that sequential decision-making poses greater challenges that better reflect real-world clinical scenarios. 3) Similarly, the improved performance observed after removing interactive information gathering (w/o Inter.) confirms the effectiveness of interactivity in our benchmark design. These results collectively suggest that while both sequentiality and interactivity make the benchmark more challenging, they are essential components that better simulate the complexity of real-world clinical decision-making processes.
Key components for MedChain-Agent: To assess the impact of individual components in our framework, we conducted ablation experiments, with results shown in Table 3. These findings highlight the crucial roles of both the Feedback mechanism and the MedCase-RAG module in enhancing overall performance. Each component independently contributes to significant improvements across various metrics. Notably, the MedCase-RAG module alone boosts Specialty Referral (Level 1) to 0.5928 and yields substantial gains in History-taking and Examination. When combined, Feedback and MedCase-RAG demonstrate a synergistic effect, achieving an impressive average score of 0.5269. This underscores the complementary nature of these components in optimizing our framework’s capabilities.
@@ 6 Conclusion
In this paper, we introduced MedChain, a novel benchmark for evaluating large language models in clinical decision-making, and MedChain-Agent, an innovative multi-agent framework designed to address the complexities of real-world medical scenarios. Our work bridges a critical gap between current AI capabilities and the nuanced realities of clinical practice by incorporating personalization, sequentiality, and interactivity into the evaluation process. MedChain-Agent’s sophisticated multi-agent system, enhanced by a feedback mechanism and the MedCase-RAG module, demonstrates unprecedented performance in sequential clinical decision-making tasks, even outperforming proprietary models when combined with open-source LLMs. By providing a more realistic and challenging evaluation framework, this research sets a new standard for evaluating and developing medical AI systems, paving the way for its responsible integration into clinical practice.
Acknowledgement
We are particularly indebted to the administrators of the iiyi website for their generosity in allowing us to utilize their data for our research purposes. We would like to acknowledge the assistance provided by Claude-3.5 in proofreading our manuscript for grammatical accuracy and in facilitating the creation of LaTeX tables.
Limitations
@@ Acknowledgement
  • Data Source Diversity: The MedChain benchmark is constructed from 12,163 electronic health records obtained from the Chinese medical website “iiYi." Although this dataset is extensive and covers 19 medical specialties and 156 sub-categories, it is derived from a single source. In our future work, we will incorporate additional data sources from different regions or healthcare systems to further enrich the dataset, providing a broader spectrum of clinical scenarios and enhancing the generalizability of the benchmark.
  • Patient Interaction Simulation: In our interactive environment, the patient responses during the history-taking stage are generated by the Gemma 2 language model. While this approach ensures consistency and control in evaluating the LLM-based agent, the real patient interactions can be more varied and complex. Future work could explore more advanced patient simulators or incorporate real dialogue data to capture a wider range of communication styles and behaviors.
@@ Abstract.
Appendix A Benchmark Construction
@@ A.1 Dataset Standardization
To ensure consistency and comparability across the benchmark, we standardized the classification of examination items into two main categories: Physical Examinations and Auxiliary Examinations. Physical Examinations include evaluations of various body systems and general health indicators, while Auxiliary Examinations encompass different imaging techniques and laboratory tests. We utilized GPT-4o to extract and classify examination items from each case, followed by manual verification to ensure accuracy. For medical imaging, we classified images into seven types, and manual review ensured the correctness of the classifications. Additionally, treatment items were extracted and categorized from each case. This standardization process ensures that the dataset is consistent, facilitating accurate and comparable evaluations of LLM performance. Figure 5 shows the statistics of different departments and Figure 4 demonstrate a case after standardization.
Figure 4: Case presentation "77_Ovarian Carcinoid with Mature Cystic Teratoma: A Case Report."

### v2

@@ 4.2 MedCase-RAG
Base LLM
GPT-4o-mini openai2023gpt4v 0.5449 0.2871 0.3399 0.5112 0.4177 0.3930 0.4156
GPT-o3 0.5495 0.2866 0.3493 - 0.4891 0.3343 0.4017
Claude-3.5-sonnet team2023gemini 0.5681 0.3050 0.3562 0.5018 0.4207 0.4053 0.4262
MedGemma sellergren2025medgemma 0.5063 0.1975 0.3315 0.6324 0.4668 0.3558 0.4105
LLaVA liu2024visual 0.3240 0.0730 0.3182 0.5165 0.3667 0.1060 0.2841
Qwen2 wang2024qwen2 0.4975 0.2215 0.4226 0.4829 0.4530 0.2193 0.3828
InternVL2 wang2024qwen2 0.4811 0.1935 0.4645 0.4490 0.4367 0.2903 0.3859
Baichuan dou2025baichuan 0.2959 0.0532 0.3264 - 0.4297 0.2591 0.2728
HuaTuoGPT zhang2023huatuogpt 0.0707 0.0207 0.3465 - 0.4163 0.1780 0.2064
MedReason wu2025medreason 0.4602 0.1258 0.3315 - 0.4741 0.2993 0.3381
FineMedLM-o1 yu2025finemedlm 0.1502 0.0124 0.3969 - 0.2772 0.1345 0.1942
Single-agent
Zero-shot 0.4811 0.1935 0.3450 0.4994 0.4572 0.2817 0.3763
Few-shot 0.5584 0.2481 0.4870 0.3758 0.4382 0.3553 0.4105
CoT wei2022chain 0.5698 0.1484 0.3750 0.6396 0.4382 0.3421 0.4189
RAG lewis2020retrieval 0.5927 0.2467 0.4732 0.6408 0.4167 0.4524 0.4704
Self-consistency wang2022self 0.5143 0.2319 0.4213 0.4144 0.4198 0.3441 0.3910
Multi-agent DyLAN liu2023dynamic 0.4415 0.1731 0.4434 0.4980 0.3863 0.2983 0.3734
AutoGen wu2023autogen 0.5228 0.2347 0.4528 0.4559 0.4250 0.3590 0.4084
MedAgent tang2023medagents 0.3830 0.2039 0.4454 0.4456 0.4102 0.3673 0.3759
MDAgent kim2024adaptive 0.2398 0.1343 0.4240 0.4983 0.3973 0.3620 0.3426
MDAgent + RAG kim2024adaptive 0.4754 0.2063 0.4412 0.5346 0.4198 0.4371 0.4190
MedChain-Agent 0.5873 0.3505 0.5836 0.6566 0.4807 0.4613 0.5200
@@ 5.1 Experimental Setup
We split the dataset into training, validation, and testing sets with a ratio of 7:1:2. Our study evaluates both single-agent and multi-agent systems. For base LLM, we test two closed-source models (gpt-4o-mini openai2023gpt4v , and claude-3.5-sonnet team2023gemini ) and four open-source models (InternVL2-8b chen2023internvl , llava-llama-3-8b-v1_1 liu2024visual , HuaTuoGPT zhang2023huatuogpt , and Qwen2-7B-Instruct wang2024qwen2 ), with model weights obtained from official Hugging Face repositories. In the single-agent evaluation, we compare with zero-shot manner, few-shot manner, CoT wei2022chain , Self-consistency wang2022self and RAG lewis2020retrieval . In the multi-agent evaluation, we compare MedChain-Agent against DyLAN liu2023dynamic , AutoGen wu2023autogen , MedAgent tang2023medagents and MDAgent kim2024adaptive . All agent framework are based on InternVL2-8b chen2023internvl . The deployment was conducted using the LMDeploy framework 2023lmdeploy . All tests executed on NVIDIA A100 GPUs featuring 80GB of memory. To enhance output stability and reliability across all experiments, we consistently set the temperature parameter to 0. The experiments were conducted in Chinese.
Table 2: Performance Comparison of LLM-based Agents across Other Diagnosis Benchmarks. Implementation of agent frameworks based on InternVL2, evaluated on multiple medical benchmarks including MedQA jin2021disease , PubMedQA jin2019pubmedqa , PathVQA he2020pathvqa , and MedBullets chen2024benchmarking .
Framework Method MedQA PubMedQA PathVQA MedBullets Average
Single-agent Zero-shot 0.426 0.668 0.449 0.490 0.508
Few-shot 0.477 0.648 0.448 0.503 0.519
CoT 0.470 0.714 0.465 0.500 0.537
Self-consistency 0.460 0.688 0.482 0.500 0.533
Multi-agent MedAgents 0.501 0.622 0.569 0.435 0.532
MDAgents 0.435 0.744 0.582 0.422 0.546
AutoGen 0.395 0.656 0.568 0.448 0.517
DyLAN 0.414 0.610 0.540 0.448 0.503
MedChain-Agent 0.462 0.746 0.621 0.474 0.576
5.2 Benchmark Performance Results
The results of our evaluation in the MedChain  are presented in Table 1. Our analysis yields two significant insights:
(1) Sequential decision-making tasks continue to pose significant challenges, even for advanced models. For instance, within the single-agent frameworks, GPT-4o-mini and InternVL2 achieve average scores of 0.4156 and 0.3859, respectively. These results indicate that despite their sophistication, these models struggle to maintain consistent performance across the sequential stages of clinical decision-making, highlighting the inherent difficulty of these tasks.
(2) The integration of the MedChain-Agent framework with open-source LLMs demonstrates significant superiority over proprietary models like GPT-4o-mini. The substantial performance gain observed with MedChain-Agent (average score of 0.5200) implies that our framework can leverage the strengths of open-source LLMs to achieve superior outcomes. This suggests that open-source models, when enhanced with our framework, are not only competitive but can also excel in handling intricate medical decision-making tasks.
5.3 Evaluation in Existing Diagnosis Datasets
Moreover, we evaluate MedChain-Agent on several well-established medical QA datasets, including MedQA jin2021disease , PubMedQA jin2019pubmedqa , PathVQA he2020pathvqa , and MedBullets chen2024benchmarking . As shown in Table 2, our framework demonstrates strong performance compared to both single-agent baselines and multi-agent alternatives. MedChain-Agent achieves the highest average score (0.576). This consistent superiority across various medical QA benchmarks further validates the effectiveness of our framework, even on simpler, more structured tasks that differ from our real-world clinical scenarios.
5.4 Ablation Studies and Discussion
@@
0.5523 0.2228 0.3285 0.6369 0.4724 0.3915 0.4341
0.5739 0.2906 0.4222 0.6377 0.4299 0.4209 0.4692
0.5928 0.3353 0.5801 0.6488 0.4699 0.4568 0.5140
0.5873 0.3505 0.5836 0.6566 0.4804 0.4613 0.5200
Table 4: Generalizability of MedChain-Agent with Various Base LLM. We apply the MedChain-Agent framework to different base LLMs to validate its generalizability. Performance changes are highlighted with light green for improvements and light red for decreases. Results demonstrate that the MedChain-Agent framework consistently brings performance gains across different base LLMs.
Base LLM Specialty Referral History-taking Examination Diagnosis Treatment Average
Level 1 Level 2
GPT-4o-mini 0.6065 (+0.0616) 0.2302 (-0.0569) 0.6205 (+0.2806) 0.6279 (+0.1167) 0.4975 (+0.0798) 0.4329 (+0.0399) 0.5026 (+0.0870)
HuaTuoGPT 0.2554 (+0.1847) 0.0097 (-0.0110) 0.3722 (+0.0257) - 0.4159 (-0.0004) 0.1450 (-0.0330) 0.2396 (+0.033)
Qwen2 0.5818 (+0.0843) 0.2781 (+0.0566) 0.5962 (+0.1736) 0.6534 (+0.1705) 0.4628 (+0.0098) 0.4315 (+0.2122) 0.5006 (+0.1178)
InternVL2 0.5873 (+0.1062) 0.3505 (+0.1570) 0.5836 (+0.1191) 0.6566 (+0.2076) 0.4806 (+0.0439) 0.4613 (+0.1710) 0.5200 (+0.1341)
Table 5: The ablation study for three key characteristics in MedChain. This table presents the impact of personalization, interactivity, and sequentiality on Diagnosis and Treatment tasks. The arrows $\uparrow\downarrow$ next to settings indicate expected performance change direction. The arrows next to results show actual changes (highlighted in gray when matching expectations).
Setting Model Diagnosis Treatment
Full MedAgent 0.4106 0.3673
MDAgent 0.3959 0.3620
gpt-4o-mini 0.4157 0.3930
InternVL2 0.4378 0.4472
MedChain-Agent 0.4802 0.4613
a) w/o Person. $\downarrow$ MedAgent 0.3075 $\downarrow$ 0.3754 $\uparrow$
@@
MedChain-Agent 0.4807 $\uparrow$ 0.4743 $\uparrow$
c) w/o Inter. $\uparrow$ MedAgent 0.3129 $\downarrow$ 0.4109 $\uparrow$
MDAgent 0.3998 $\uparrow$ 0.3627 $\uparrow$
@@
InternVL2 0.4550 $\uparrow$ 0.3173 $\downarrow$
MedChain-Agent 0.4634 $\downarrow$ 0.5207 $\uparrow$
(1) Ablation Study for Key components in MedChain-Agent: To assess our framework’s components, we conduct ablation studies as shown in Table 3. Both the Feedback mechanism and MedCase-RAG module demonstrate significant individual contributions to overall performance. The Feedback mechanism alone improves the average score from 0.4341 to 0.4692, with notable gains in History-taking and Level 2 Specialty referral. MedCase-RAG shows stronger individual impact, boosting the average score to 0.5140, with substantial improvements in History-taking and Specialty referral tasks. While MedCase-RAG excels in diagnostic phases, the Feedback mechanism appears more beneficial for Treatment tasks. When combined, these components show synergistic effects, achieving the highest average performance (0.5200) and optimal scores across five of six evaluated tasks, validating their complementary roles in enhancing clinical reasoning capabilities.
(2) Ablation Study for Three Key Characteristics in MedChain: To validate the effectiveness of personalization, interactivity, and sequentiality within our benchmark, we conduct an ablation study as shown in Table 5. We systematically remove each characteristic and observe its impact on model performance across Diagnosis and Treatment tasks. ’w/o Person’ means all detailed patient profiles are omitted from the input, resulting in a lack of personalized information that makes correct diagnosis more challenging for the model. ’w/o Seq’ means we use the ground truth from the previous stage as input to the next stage, rather than using the model’s previous output, which simplifies the benchmark. ’w/o Inter’ means we directly provide all patient examination results as input without requiring the agent to autonomously inquire about the patient’s condition, also simplifying the benchmark. Removing patient-specific information (w/o Person.) consistently degrades performance across all models in diagnosis tasks (with drops ranging from 2.51% to 10.31%), demonstrating that personalized information is crucial for accurate clinical decision-making. Interestingly, when removing the sequential dependency between stages (w/o Seq.), most models show improved performance, indicating that sequential decision-making poses greater challenges that better reflect real-world clinical scenarios. Similarly, the improved performance observed after removing interactive information gathering (w/o Inter.) confirms the effectiveness of interactivity in our benchmark design. These results collectively suggest that while both sequentiality and interactivity make the benchmark more challenging, they are essential components that better simulate the complexity of real-world clinical decision-making processes.
(3) Generalizability for MedChain-Agent: To evaluate the generalizability of our framework, we apply MedChain-Agent to various base LLMs as shown in Table 4. The results demonstrate substantial performance improvements across most models and tasks. InternVL2 shows the most significant enhancement with a 13.41% average improvement, followed by Qwen2 (11.78%) and GPT-4o-mini (8.70%). Even with HuaTuoGPT, which has relatively lower baseline performance, our framework still achieves a 3.33% improvement. Notably, the improvements are particularly pronounced in complex tasks such as History-taking and Examination, where gains of up to 28.06% are observed. These consistent enhancements across diverse models and tasks strongly validate the framework’s robust generalizability and effectiveness in medical decision-making scenarios.
@@ 6 Conclusion
In this paper, we introduced MedChain, a novel benchmark for evaluating LLM-based agents in clinical decision-making that authentically reflects real-world medical practice through three essential characteristics: personalization, interactivity, and sequentiality. Our comprehensive dataset encompasses 12,163 diverse clinical cases across 19 medical specialties, including 7,338 medical images with corresponding reports, providing a robust foundation for evaluating AI systems in complex healthcare scenarios. To address the challenges presented by this benchmark, we also introduced MedChain-Agent, a multi-agent framework enhanced by feedback mechanisms and MedCase-RAG, which demonstrates superior performance across various clinical tasks. This work establishes new benchmarks for evaluating medical AI systems and provides practical solutions for enhancing their clinical decision-making capabilities. As AI continues to evolve in healthcare applications, frameworks that can navigate the complexities of real-world clinical scenarios will be increasingly valuable for improving patient care and supporting healthcare professionals.
7 Limitations
@@
1) Data Source Diversity: The MedChain benchmark is constructed from 12,163 electronic health records obtained from the Chinese medical website “iiYi." Although this dataset is extensive and covers 19 medical specialties and 156 sub-categories, it is derived from a single source. Additionally, there exists a notable imbalance in data distribution across different medical specialties, with some departments having significantly more cases than others. This imbalance may introduce bias in model evaluation and limit performance in underrepresented specialties. In our future work, we will incorporate additional data sources from different regions or healthcare systems to further enrich the dataset and address the specialty imbalance.
2) Patient Interaction Simulation: In our interactive environment, the patient responses during the history-taking stage are generated by the Gemma 2 language model. While this approach ensures consistency and control in evaluating the LLM-based agent, the real patient interactions can be more varied and complex. Future work could explore more advanced patient simulators or incorporate real dialogue data to capture a wider range of communication styles and behaviors.
@@ Abstract.
Appendix D discuss the difference between MedChain and several similar related works.
Appendix A Benchmark Construction and Evaluation
@@ A.1 Dataset Standardization
To ensure consistency and comparability across the benchmark, we standardized the classification of examination items into two main categories: Physical Examinations and Auxiliary Examinations. Physical Examinations include evaluations of various body systems and general health indicators, while Auxiliary Examinations encompass different imaging techniques and laboratory tests. We utilized GPT-4o to extract and classify examination items from each case, followed by manual verification to ensure accuracy. For medical imaging, we classified images into seven types, and manual review ensured the correctness of the classifications. Additionally, treatment items were extracted and categorized from each case. This standardization process ensures that the dataset is consistent, facilitating accurate and comparable evaluations of LLM performance. Figure 4 and Figure 5 demonstrate a case after standardization in English and Chinese. Figure 6 shows the statistics of different departments.
Figure 4: Case Report "77_Ovarian Carcinoid with Mature Cystic Teratoma: A Case Report."
Figure 5: Case Report Chinese Version with Corresponding Medical Imaging. "77_Ovarian Carcinoid with Mature Cystic Teratoma: A Case Report." (a), (b) and (c) Medical Imaging. (d) Chinese version.